OpenStudy (anonymous):

Find the solution of this system of equations. -2x-4y=42 2x-8y=18

OpenStudy (anonymous):

lets say we take your first equation and subtract the second equations from it

OpenStudy (anonymous):

that being -2x-4y=42 2x-8y=18 -2+2x-4y+(-8y) = 42 +18

OpenStudy (anonymous):

-2x+2x-4y+(-8y) = 42 +18 (CORRECTED) and I also meant to say add both the equations together. sorry for the mix up

OpenStudy (anonymous):

so therefore -2x+2x-4y+(-8y) = 42 +18 -12y = 60 y = 60/ -12 y = -5 sub y = -5 into 2x-8y=18 2x - 8(-5) = 18 2x + 40 =18 2x = 18-40 x = -11